# Retrieve Anyone: A General-purpose Person Re-identification Task with Instructions

### v1

@@ Abstract
Human intelligence can retrieve any person according to both visual and language descriptions. However, the current computer vision community studies specific person re-identification (ReID) tasks in different scenarios separately, which limits the applications in the real world. This paper strives to resolve this problem by proposing a new instruct-ReID task that requires the model to retrieve images according to the given image or language instructions. Our instruct-ReID is a more general ReID setting, where existing ReID tasks can be viewed as special cases by designing different instructions. We propose a large-scale OmniReID benchmark and an adaptive triplet loss as a baseline method to facilitate research in this new setting. Experimental results show that the baseline model trained on our OmniReID benchmark can improve +0.5%, +3.3% mAP on Market1501 and CUHK03 for traditional ReID, +2.1%, +0.2%, +15.3% mAP on PRCC, VC-Clothes, LTCC for clothes-changing ReID, +12.5% mAP on COCAS+ real2 for clothestemplate based clothes-changing ReID when using only RGB images, +25.5% mAP on COCAS+ real2 for our newly defined language-instructed ReID. The dataset, model, and code will be available at https://github.com/hwz-zju/Instruct-ReID.
@@ 1 Introduction
Identifying individuals exhibiting significant appearance variations from multimodal descriptions is a fundamental aspect of human intelligence with broad applications [29; 27]. To imbue our machine with this capability, person re-identification (ReID) [52; 50] has been introduced to retrieve images of the target person from a vast repository of surveillance videos or images across locations and time [9]. In recent years, significant advancements have been made in developing precise and efficient ReID algorithms and establishing benchmarks covering various scenarios, such as traditional ReID [49; 4; 40], clothes-changing ReID (CC-ReID) [16; 11; 17; 35; 13], clothes template based clothes-changing ReID (CTCC-ReID) [46; 23], and text-to-image ReID (T2I-ReID) [3; 51; 24; 45]. However, focusing solely on one specific scenario, specialized ReID models possess inherent limitations. For instance, customers must employ distinct models to retrieve persons according to languages or clothes templates. To facilitate real-world deployment, there is a pressing need to devise one generic framework capable of re-identifying individuals across all scenarios mentioned above.
This paper proposes a new instruct-ReID task where existing ReID settings can be formulated as special cases. Specifically, the instruct-ReID uses query images and multimodal instructions as the model inputs and requires the model to retrieve the images of the same person in the query image from the gallery following the instructions. Using language or image instructions, models trained on the instructed-ReID task can be specialized to tackle diverse ReID tasks (Fig. 1a). For example, the clothes-changing ReID can be viewed as using the instruction "Ignore clothes" to retrieve images. As another example, clothes template based clothes-changing ReID can utilize a clothes template image as instruction. The proposed instruct-ReID task offers three significant advantages: easy deployment, improved performance, and convenient extension to new ReID tasks. First, it enables cost-effective and convenient deployment in real-world applications. Unlike existing ReID approaches limited to specific tasks, instruct-ReID allows for utilizing a single model for various ReID scenarios, which is more practical in real applications. Second, as all existing ReID tasks can be considered special cases of instruct-ReID, we can unify the training sets of these tasks to exploit the benefits of more data and diverse annotations across different tasks – leading to enhanced performance. Third, instruct-ReID can address a new language-instructed ReID task which requires the model to retrieve persons following language instructions. This setting is very practical in real applications because it enables the customers to retrieve images they are particularly interested in through language descriptions. For example, customers can retrieve a woman wearing a black coat using one of her pictures and language instructions "Wear black coat".
@@
Figure 1: (a) We proposed a new instruct-ReID task that unites various ReID tasks. Traditional ReID: The instruction is "Do not change clothes". The model is expected to retrieve images of the same person without altering image attributes, such as clothing. Clothes-changing ReID: The instruction is "Ignore clothes". The model should retrieve images of the same person even when wearing different outfits. Clothes template based clothes-changing ReID: The instruction is a cropped clothes image. The model should retrieve images of the same person wearing the provided clothing. Language-instructed ReID: The instruction is several sentences describing pedestrian attributes. The model is required to retrieve images of the same person described in the provided sentences. (b) Our proposed method improves the performance of various person ReID tasks by a considerable margin.
@@
To facilitate research in instruct-ReID, we introduce a new benchmark called OmniReID derived from 8 datasets of 4 different ReID tasks. The OmniReID benchmark exhibits two appealing characteristics. First, it emphasizes diversity by incorporating images from various domains, including surveillance scenarios and movie clips. The diversity ensures that the trained models are robust and can effectively handle ReID tasks in various real-world scenarios. Second, the OmniReID benchmark achieves comprehensiveness by offering evaluation datasets to assess various ReID tasks comprehensively, which facilitates evaluating the generalization ability of diverse ReID methods.
@@
Based on OmniReID, we design a baseline method with the proposed adaptive triplet loss for instruct-ReID. The typical triplet loss [34] only defines positive/negative pairs by identities, failing to align with instructions. Therefore, we propose a novel adaptive triplet loss to learn a metric space that preserves identity and instruction similarities. Specifically, we design an adaptive margin between two query-instruction pairs based on instruction similarities to pull features with similar instructions close and push features with different instructions apart. This loss incorporates the instruction information into the features representation and can optimize the model to learn a metric space where similar instructions lead to closer features of query-instance pairs.
@@
In summary, the contributions of this paper are three folds. (1) We propose a new instruct-ReID task, where existing traditional ReID, clothes-changing ReID, clothes template based clothes-changing ReID, and language-instructed ReID can be viewed as special cases. (2) To facilitate research on instruct-ReID, we establish a large-scale and comprehensive OmniReID benchmark consisting 8 publicly available datasets. (3) We propose an adaptive triplet loss to supervise the feature distance of two query-instruction pairs to consider identity and instruction alignments. Our method achieves state-of-the-art results on 7 datasets throughout 4 tasks and on-par results on 1 dataset on traditional ReID tasks. Specifically, our method improves +3.3%, +0.5% mAP on CUHK03, Market1501 for traditional ReID, +2.1%, 15.3% mAP on PRCC, LTCC for clothes-changing ReID when using RGB images only, +12.5% mAP on COCAS+ real2 for clothes template based clothes changing ReID, +25.5% mAP on COCAS+ real2 for our newly defined language-instructed ReID.
@@ 2 Related work
Person Re-identification. Person re-identification aims to determine the target images of the same person in the gallery set given the query image. To support the ReID task on all-weather, various tasks are conducted on the scenarios with changing environments, perspectives, and poses [44; 4; 50; 3; 51]. Traditional ReID mainly focuses on dealing with indoor/outdoor and camera pose problems when the target person wears the same clothes. Recently, to extend the application scenarios, clothes-changing ReID (CC-ReID) [46] and clothes template based clothes-changing ReID (CTCC-ReID) [23] are proposed. While CC-ReID forces the model to learn clothe-invariant features [16; 11], CCTC-ReID further extracts clothes template features [23] to retrieve the image of the person wearing the same clothes. In this paper, we introduce a new Instruct-ReID task, which can be viewed as a superset of the existing ReID tasks by incorporating instruction information into identification.
Instruction Tuning. Instruction Tuning was first proposed in the natural language processing (NLP) community to enable large language models (LLM) to execute specific tasks by following natural language instructions. LLM instruction-tuned models, e.g., FLAN-T5 [6], InstuctGPT [31]/ChatGPT [30], can effectively prompt the ability on zero- and few-shot transfer tasks. A few works have borrowed the idea from language to vision. Flamingo [1], BLIP-2 [20], and KOSMOS-1 [15] learning with image-text pairs also show promising generalization on visual understanding tasks. While these LLM-based methods aim to generate convincing language responses following the image or language instructions, we focus on retrieving the correct person following the given instructions by tuning a vision transformer.
@@ 3 OmniReID Benchmark
To facilitate research on Instruct-ReID, we propose the OmniReID benchmark including a large-scale pretraining dataset based on 8 publicly available datasets with visual and language annotations. The comparison with the existing ReID benchmark is illustrated in Tab. 1.
@@
MSMT17 [40] 30248 1041 indoor/outdoor
CUHK03 [25] 7365 767 indoor
Market1501 [48] 12936 751 outdoor
PRCC [43] 17896 150 indoor
VC-Clothes [39] 9449 256 synthetic
COCAS+ Real1 [23] 34469 2800 indoor/outdoor
OmniReID 121786 5842 indoor/outdoor/synthetic
@@
Protocols. To enable all-purpose person ReID, we collect massive public datasets from various domains and use their training subset as our training subset, including Market1501 [48], MSMT17 [40], CUHK03 [25] for traditional ReID, PRCC [43], VC-Clothes [39], LTCC [32] for clothes-changing ReID, COCAS+ Real1 [23] for clothes template based clothes-changing ReID and language-instructed ReID, forming 121,786 images and 5,842 identities. To fairly compare our method with state-of-the-art methods, the trained models are evaluated on LTCC, PRCC, VC-Clthoes, COCAS+ Real2, Market1501, MSMT17, and CUHK03 test subsets without finetuning. We search the target images with query images and the instructions. Since a query image has multiple target images in the gallery set and CMC (Cumulative Matching Characteristic) curve only reflect the retrieval precision of most similar target images, we also adopt mAP (mean Average Precision) to reflect the overall ranking performance w.r.t. all target images.
@@
Language Annotation Generation. Similar to Text-to-Image ReID datasets, e.g., CUHK-PEDES [24], language annotations in OmniReID are several sentences that describe the visual appearance of pedestrians. We divide our annotation process into pedestrian attribute generation and attribution-to-language transformation.
@@
Pedestrian Attribute Generation. To obtain a comprehensive and varied description of an individual, we employ an extensive collection of attribute words describing a wide range of human visual characteristics. The collection contains 20 descriptions, including full-body clothing, hair color, hairstyle, gender, age, actions, posture, and accessories such as umbrellas or satchels. Professional annotators manually label all the pedestrian attributes. We provide a practical illustration with the attribute collection in Fig. 2(a). By utilizing the instructions on the well-defined attribute combination, models can further enhance their ability to identify the target person. All detailed pedestrian attributes are provided in the supplementary materials.
@@
Attribute-to-Language Transformation. Compared with discrete attribute words, language is more natural for consumers. To this end, we transform these attributes into multiple sentences using the Alpaca-LoRA large language model. Specifically, we ask the Alpaca-LoRA with the following sentences: "Generate sentences to describe a person. The above sentences should contain all the attribute information I gave you in the following." Annotators carefully check generated annotations to ensure the correctness of the language instructions. We provide more language annotations of images in the supplementary materials.
@@
Figure 2: (a) We first generate attributes for a person and then transform attributes into sentences by a large language model. (b) We crop the upper clothes and search them online for clothes templates.
@@
Visual Annotation Generation. Visual annotations are generally images that describe the characteristics of pedestrians. In this paper, we select clothes as the visual annotations because they are viewed as the most significant visual characteristics of pedestrians. To obtain high-quality visual annotations, we first crop the upper clothes from the source images (Clothes Copping) and search for the upper clothes on the internet to get the most corresponding cloth-template images (Clothes-template Crawling) as visual annotations. Since each person wears the same clothes in the traditional ReID dataset, we only annotate the clothes-changing ReID datasets where each person wears multiple clothes to ease the burden of annotations. We detail the process in Fig. 2 (b).
@@
Clothes Cropping. We use a human parsing model, SCHP [22], to generate the segmentation mask of the upper clothes and then crop the corresponding rectangle upper clothes patches from the original images. These bounding boxes of upper clothes are then manually validated.
@@
Figure 3: The overall architecture of the proposed method. The instruction is fed into the instruction encoder to extract instruction features. The features are then propagated into the editing transformer to capture instruction-edited features. We exploit adaptive triplet loss and identification loss to train the network. For the testing stage, we use the CLS token for retrieval.
4 RAM: Retrieve Anyone Model
The proposed retrieve anyone model is shown in Fig. 3. Our model consists of three parts: instruction generation for different ReID tasks, an editing transformer $\mathcal{E}_{e}$ and an instruction encoder $\mathcal{E}_{i}$ which is the fixed FashionCLIP model [5], which we show better discriminative abilities than CLIP model [33] in the supplementary materials. Given an instruction $\mathbf{T}$ associated with a query image $\mathbf{I}$ (Sec. 4.1), we extract instruction features $\mathbf{F}_{\mathbf{T}}$ by the instruction encoder $\mathcal{E}_{i}$. Then the extracted instruction feature $\mathbf{F}_{\mathbf{T}}$ and the query image tokens are fed into the designed editing transformer to get the features $\mathbf{F}$ edited by institutions (Sec. 4.2). We impose an identification loss $\mathcal{L}_{id}$, which is the classification loss on identities, and an adaptive triplet loss $\mathcal{L}_{atri}$ on $\mathbf{F}$ to supervise the learning of model. The adaptive triplet loss is utilized to pull features with the same identity and similar instructions close and push features with dissimilar instructions apart (Sec. 4.3).

### v4

# Instruct-ReID: A Multi-purpose Person Re-identification Task with Instructions

@@ Abstract
Human intelligence can retrieve any person according to both visual and language descriptions. However, the current computer vision community studies specific person re-identification (ReID) tasks in different scenarios separately, which limits the applications in the real world. This paper strives to resolve this problem by proposing a new instruct-ReID task that requires the model to retrieve images according to the given image or language instructions. Our instruct-ReID is a more general ReID setting, where existing 6 ReID tasks can be viewed as special cases by designing different instructions. We propose a large-scale OmniReID benchmark and an adaptive triplet loss as a baseline method to facilitate research in this new setting. Experimental results show that the proposed multi-purpose ReID model, trained on our OmniReID benchmark without fine-tuning, can improve +0.5%, +0.6%, +7.7% mAP on Market1501, MSMT17, CUHK03 for traditional ReID, +6.4%, +7.1%, +11.2% mAP on PRCC, VC-Clothes, LTCC for clothes-changing ReID, +11.7% mAP on COCAS+ real2 for clothes template based clothes-changing ReID when using only RGB images, +24.9% mAP on COCAS+ real2 for our newly defined language-instructed ReID, +4.3% on LLCM for visible-infrared ReID, +2.6% on CUHK-PEDES for text-to-image ReID. The datasets, the model, and code will be available at https://github.com/hwz-zju/Instruct-ReID.
@@ 1 Introduction
Identifying individuals exhibiting significant appearance variations from multimodal descriptions is a fundamental aspect of human intelligence with broad applications [34, 31]. To imbue our machine with this capability, person re-identification (ReID) [67, 66] has been introduced to retrieve images of the target person from a vast repository of surveillance videos or images across locations and time [10]. In recent years, significant advancements have been made in developing precise and efficient ReID algorithms and establishing benchmarks covering various scenarios, such as traditional ReID [65, 6, 47], clothes-changing ReID (CC-ReID) [18, 12, 20, 41, 15], clothes template based clothes-changing ReID (CTCC-ReID) [57, 28], visible-infrared ReID (VI-ReID) [32, 50, 51, 59] and text-to-image ReID (T2I-ReID) [3, 4, 68, 26, 56]. However, focusing solely on one specific scenario, specialized ReID models possess inherent limitations. For instance, customers must deploy distinct models to retrieve persons according to the query information, which significantly increases the cost for model training and results in inconvenience for the application. To facilitate real-world deployment, there is a pressing need to devise one generic framework capable of re-identifying individuals across all scenarios mentioned above.
@@
Figure 1: (a) We proposed a new instruct-ReID task that unites various ReID tasks. Traditional ReID: The instruction may be “Do not change clothes”. Clothes-changing ReID: The instruction may be “Ignore clothes”. Clothes template based clothes-changing ReID: The instruction is a cropped clothes image and the model should retrieve the same person wearing the provided clothing. Language-instructed ReID: The instruction is several sentences describing pedestrian attributes. The model is required to retrieve the person described by the instruction. Visible-Infrared ReID: The instruction can be “Cross modality”. Text-to-image ReID: The model retrieves images according to the description sentence. (b) Our proposed method improves the performance of various person ReID tasks by a unified retrieval model.
This paper proposes a new multi-purpose instruct-ReID task where existing 6 ReID settings can be formulated as its special cases. Specifically, the instruct-ReID uses query images and multimodal instructions as the model inputs, requiring the model to retrieve the same identity images from the gallery following the instructions. Using language or image instructions, models trained on the instruct-ReID task can be specialized to tackle diverse ReID tasks (Fig. 1a). For example, the clothes-changing ReID can be viewed as using the instruction “Ignore clothes” to retrieve. As another example, clothes template-based clothes-changing ReID can utilize a clothes template image as instruction. The proposed instruct-ReID task offers three significant advantages: easy deployment, improved performance, and easy extension to new ReID tasks. First, it enables cost-effective and convenient deployment in real-world applications. Unlike existing ReID approaches limited to specific tasks, instruct-ReID allows for utilizing a single model for various ReID scenarios, which is more practical in real applications. Second, as all existing ReID tasks can be considered special cases of instruct-ReID, we can unify the training sets of these tasks to exploit the benefits of more data and diverse annotations across various tasks – leading to enhanced performance. Third, instruct-ReID addresses a new language-instructed ReID task which requires the model to retrieve persons following language instructions. This setting is very practical in real applications for it enables customers to retrieve images they are particularly interested in through language descriptions. For example, customers can retrieve a woman wearing a black coat using one of her pictures and language instructions “Wear black coat”.
@@
To facilitate research in instruct-ReID, we introduce a new benchmark called OmniReID derived from 12 datasets  [^1] The discussion on ethical risks is provided in supplementary materials. of 6 different ReID tasks. The OmniReID benchmark exhibits two appealing characteristics. First, it emphasizes diversity by incorporating images from various domains, including surveillance scenarios and movie clips. The diversity ensures that the trained models are robust and can effectively handle ReID tasks in various real-world scenarios. Second, the OmniReID benchmark achieves comprehensiveness by offering evaluation datasets to assess various ReID tasks comprehensively, which facilitates evaluating the generalization ability of diverse ReID methods.
@@
Based on OmniReID, we design a method with the proposed adaptive triplet loss for instruct-ReID. The typical triplet loss [39] only defines positive/negative pairs by identities, failing to align with instructions. Therefore, we propose a novel adaptive triplet loss to learn a metric space that preserves identity and instruction similarities. Specifically, we design an adaptive margin between two query-instruction pairs based on instruction similarities to pull features with similar instructions close and push features with different instructions apart. This loss incorporates the instruction information into the features representation and optimizes the model to learn a metric space where similar instructions lead to closer features of query-instance pairs.
@@
In summary, the contributions of this paper are three folds. (1) We propose a new instruct-ReID task, where existing traditional ReID, clothes-changing ReID, clothes template based clothes-changing ReID, visible-infrared ReID, text-to-image ReID and language-instructed ReID can be viewed as special cases. (2) To facilitate research on instruct-ReID, we establish a large-scale and comprehensive OmniReID benchmark consisting 12 publicly available datasets. (3) We propose an adaptive triplet loss to supervise the feature distance of two query-instruction pairs to consider identity and instruction alignments. Our method consistently improves previous models on 10 datasets of 6 ReID tasks using ImageNet pretraining or ReID-specific pretraining [69]. For example, when using ReID-specific pretraining, our method improves +7.7%, +0.6%, +0.5% mAP on CUHK03, MSMT17, Market1501 for traditional ReID, +6.4%, +11.2%, +7.1% mAP on PRCC, LTCC, VC-Clothes for clothes-changing ReID when using RGB images only, +11.7% mAP on COCAS+ real2 for clothes template based clothes changing ReID, +4.3% mAP on LLCM for visible-infrared ReID, +2.6% mAP on CUHK-PEDES for text-to-image ReID, +24.9% mAP on COCAS+ real2 for the new language-instructed ReID.
@@ 2 Related work
Person Re-identification. Person re-identification aims to retrieve the same images of the same identity with the given query from the gallery set. To support the ReID task on all-weather application, various tasks are conducted on the scenarios with changing environments, perspectives, and poses [55, 6, 66, 4, 68]. Traditional ReID mainly focuses on dealing with indoor/outdoor problems when the target person wears the same clothes. Recently, to extend the application scenarios, clothes-changing ReID (CC-ReID) [57] and clothes template based clothes-changing ReID (CTCC-ReID) [28] are proposed. While CC-ReID forces the model to learn clothes-invariant features [18, 12], CCTC-ReID further extracts clothes template features [28] to retrieve the image of the person wearing template clothes. To capture person’s information under low-light environments, visible-infrared person ReID (VI-ReID) methods [59, 50, 51, 59] retrieve the visible (infrared) images according to the corresponding infrared (visible) images. In the absence of the query image, GNA-RNN [26] introduced the text-to-image ReID (T2I-ReID) task, which aims at retrieving the person from the textual description. However, existing tasks and methods focus on a single scenario, making it difficult to address the demands of cross-scenario tasks. In this paper, we introduce a new Instruct-ReID task, which can be viewed as a superset of the existing ReID tasks by incorporating instruction information into identification.
Instruction Tuning. Instruction Tuning was first proposed in natural language processing (NLP) to enable large language models (LLM) to execute specific tasks by following natural language instructions. LLM instruction-tuned models, e.g., FLAN-T5 [7], InstuctGPT [36]/ChatGPT [35], can effectively prompt the ability on zero- and few-shot transfer tasks. A few works have borrowed the idea from language to vision. Flamingo [2], BLIP-2 [23], and KOSMOS-1 [17] learning with image-text pairs also show promising generalization on visual understanding tasks. While these LLM-based methods aim to generate convincing language responses following the image or language instructions, we focus on retrieving the correct person following the given instructions by tuning a vision transformer.
@@ 3 OmniReID Benchmark
To facilitate research on Instruct-ReID, we propose the OmniReID benchmark including a large-scale pretraining dataset based on 12 publicly available datasets with visual and language annotations. The comparison with the existing ReID benchmark is illustrated in Tab. 1.
@@
MSMT17 [47] 30,248 1,041 indoor/outdoor
Market1501 [64] 12,936 751 outdoor
PRCC [52] 17,896 150 indoor
COCAS+ Real1 [28] 34,469 2,800 indoor/outdoor
LLCM [61] 30,921 713 indoor/outdoor
LaST [42] 71,248 5,000 indoor/outdoor
MALS [53] 1,510,330 1,510,330 synthesis
LUPerson-T [40] 957,606 - indoor/outdoor
OmniReID 4,973,044 328,604 indoor/outdoor/synthetic
@@
Protocols. To enable all-purpose person ReID, we collect massive public datasets from various domains and use their training subset as our training subset, including Market1501 [64], MSMT17 [47], CUHK03 [30] for traditional ReID, PRCC [52], VC-Clothes [46], LTCC [37] for clothes-changing ReID, LLCM [61] for visible-infrared ReID, CUHK-PEDES [26], SYNTH-PEDES [70] for text-to-image ReID, COCAS+ Real1 [61] for clothes template based clothes-changing ReID and language-instructed ReID, forming 4,973,044 images and 328,604 identities. To fairly compare our method with state-of-the-art methods, the trained models are evaluated on LTCC, PRCC, VC-Clthoes, COCAS+ Real2, LLCM, CUHK-PEDES, Market1501, MSMT17, and CUHK03 test subsets without finetuning. We search the target images with query images and the instructions. Since a query image has multiple target images in the gallery set and CMC (Cumulative Matching Characteristic) curve only reflects the retrieval precision of most similar target images, we also adopt mAP (mean Average Precision) to reflect the overall ranking performance w.r.t. all target images. We present all dataset statistics of OmniReID in the supplementary materials.
@@
Language Annotation Generation. To generate language instruction for language-instructed ReID, we annotate COCAS+ Real1 and COCAS+ Real2 with language description labels. Similar to Text-to-Image ReID datasets, e.g., CUHK-PEDES [26], language annotations in OmniReID are several sentences that describe the visual appearance of pedestrians. We divide our annotation process into pedestrian attribute generation and attribution-to-language transformation.
@@
Pedestrian Attribute Generation. To obtain a comprehensive and varied description of an individual, we employ an extensive collection of attribute words describing a wide range of human visual characteristics. The collection contains 20 attributes and 92 specific representation words, including full-body clothing, hair color, hairstyle, gender, posture, and accessories such as umbrellas or satchels. Professional annotators manually label all the pedestrian attributes. We provide a practical illustration with the attribute collection in Fig. 2(a). By utilizing instructions on the well-defined attribute combination, models can further enhance their ability to identify the target person.
@@
Attribute-to-Language Transformation. Compared with discrete attribute words, language is more natural for consumers. To this end, we transform these attributes into multiple sentences using the Alpaca-LoRA [60] large language model. Specifically, we ask the Alpaca-LoRA with the following sentences: “Generate sentences to describe a person. The above sentences should contain all the attribute information I gave you in the following.” The generated annotations are carefully checked and corrected manually to ensure the correctness of the language instructions. All detailed pedestrian attributes and more language annotations are presented in the supplementary materials.
@@
Figure 2: (a) We generate attributes for a person and then transform attributes into sentences by a large language model. (b) We crop upper clothes and search them online for clothes templates.
Figure 3: The overall architecture of the proposed method. The instruction is fed into the instruction encoder to extract instruction features (a). The features are then propagated into the editing transformer (b) to capture instruction-edited features. We exploit adaptive triplet loss and identification loss to train the network. For the testing stage, we use the CLS token for retrieval.
@@
Visual Annotation Generation. Visual annotations are images that describe the characteristics of pedestrians. In this paper, we select clothes as the visual annotations because they are viewed as the most significant visual characteristics of pedestrians. To get high-quality visual annotations, we first crop the upper clothes from the source images (Clothes Copping) and search on the internet to get the most corresponding clothes-template images (Clothes-template Crawling) as visual annotations. Since each person wears the same clothes in traditional ReID datasets, we annotate the clothes-changing LTCC dataset where each person wears multiple clothes to ease the burden of annotations. Fig. 2 (b) shows the detailed process.
@@
Clothes Cropping. We use a human parsing model SCHP [25] to generate the segmentation mask of the upper clothes and then crop the corresponding rectangle upper clothes patches from the original images. These bounding boxes of upper clothes are then manually validated.
@@
4 IRM: Instruct ReID Model
The proposed Instruct ReID model is shown in Fig. 3 (a). Our model consists of four parts: instruction generation (Sec. 4.1) for different ReID tasks, an editing transformer $\mathcal{E}_{e}$ (Sec. 4.2) and an instruction encoder $\mathcal{E}_{i}$ which is a visual language model, and a cross-model attention module $\mathcal{E}_{f}$. Given an instruction $\mathbf{T}$ associated with a query image $\mathbf{I}$, we obtain instruction features $\mathbf{F}_{\mathbf{T}}$ using the instruction encoder $\mathcal{E}i$. These extracted instruction features $\mathbf{F}_{T}$, along with query image tokens, are fed into the designed editing transformer $\mathcal{E}_{e}$ to obtain features $\mathbf{F}$ edited by instructions. Furthermore, we introduce an attention module $\mathcal{E}_{f}$ to efficiently combine features of the query image and instruction through the attention mechanism in the stacked transformer layers. Finally, the overall loss function is introduced (Sec. 4.4) for all ReID tasks, which includes our newly proposed adaptive triplet loss $\mathcal{L}_{atri}$ (Sec. 4.3) instead of traditional triplet loss.